Calculate the specific weight, specific mass, specific volume and specific gravity of a liquid having a volume of 5 cubic meters and its weight is 35 kilo newton

1
Expert's answer
2022-02-23T12:11:30-0500

Answer

Weight of liquid

"W=35KN"

So it's mass becomes

"M=\frac{W}{g}\\=\frac{35000}{9.81}\\=3567.8Kg"

Volume

"V=5m^3"

So

"Specific \space weight
\\\gamma=\frac{W}{V}\\=\frac{35000}{5}\\=7000N/m^3"

Specific mass

M=3567.8Kg

Density

"\rho =\frac{M}{V}\\=\frac{3567.8}{5}\\=711.54Kg/m^3"

Specific volume

"v=\frac{V}{M}\\=\frac{5}{3567.8}\\=0.0014m^3/kg"


Specific gravity

"Sg=\frac{\gamma}{\gamma_{at 4°C}}\\=\frac{7.000}{9.81}\\=0.71"